# Ctrl-Crash: Controllable Diffusion for Realistic Car Crashes

### v1

@@ 3.3 Conditioning Signals
Initial Image Frame (Contextual Conditioning): The initial image provides the visual grounding for generation, capturing the appearance, layout, and environment of the driving scene before any dynamic event unfolds. It is encoded using a pretrained visual encoder (VAE) and fed into the base diffusion model as context. The initial image strongly influences scene appearance and plausibility, and serves as the starting point for generating temporally coherent crash evolutions. Bounding Box Conditioning (Spatial Control): Road users are represented with 2D bounding boxes rasterized into RGB control frames. These provide spatial guidance regarding each object’s motion, size, and location over time. Each bounding box is color-coded to encode both its unique track ID (via fill color) and its object class (via border color), enabling the model to distinguish agents across frames. These control frames are encoded using the same pretrained VAE as the initial image and processed by a ControlNet and injected into the denoising process, directing the motion of agents in the generated video. See Appendix E for visualizations and details of the encoding scheme. Crash Type Conditioning (Semantic Control): Crash types are represented by a discrete class label from five categories: (0) none, (1) ego-only, (2) ego/vehicle, (3) vehicle-only, and (4) vehicle/vehicle. These indicate which agents are involved in the crash, with, for example, “vehicle/vehicle” describing a crash between two non-ego agents. The crash type index is embedded and projected into the cross-attention layers of the ControlNet encoder, allowing the model to generate outcomes consistent with high-level semantic intent.
@@ 3.4 Training Pipeline
Our training strategy is divided into two phases. In the first stage, we fine-tune the Stable Video Diffusion (SVD) model on crash-heavy video data from our preprocessed version of the MM-AU dataset (see Section 3.2). The setup is an image-to-video task with an MSE loss in latent space. This improves the model’s ability to generate plausible driving videos, including crashes, though without explicit controllability. In the second stage, we freeze the fine-tuned base model and train a ControlNet adapter module to introduce conditional controllability. This stage allows us to direct the generation using spatial and semantic control signals—namely, bounding box frames and crash type identifiers. More training and implementation details are provided in Appendix G.
Conditioning Masking. To promote robustness and controllability at inference time, we apply randomized masking of conditioning signals during training. This enables the model to generalize across varying levels of supervision and supports flexible classifier-free guidance at inference. For bounding box conditioning, we introduce a temporal dropout strategy: at each training step, a timestep $k\in[0,T]$ is sampled uniformly, and all bounding box frames from timestep $k$ onward are masked using a learnable null embedding. Unlike zero-valued tensors, the null embedding prevents the model from misinterpreting masked frames as scenes with no agents present. This teaches the model to perform plausibly even when only partial agent trajectory information is available. We adopt a curriculum schedule: bounding box masking is applied with 50% probability for the first 21k training steps, and with 100% probability thereafter (up to 31k steps), encouraging early learning from dense supervision before transitioning to partial conditioning. For semantic signals (crash type and initial image), we independently mask in the following way: with 10% probability, only the crash type is masked; with 10% probability, only the initial image is masked; and with 10% probability: both are masked. This helps prevent the model from collapsing onto any single conditioning signal and allows classifier-free guidance to function reliably across different control configurations.
3.5 Classifier-Free Guidance with Multi-Condition Factorization
Our model supports three conditioning modalities: the initial image $c_{I}$, the bounding box frames $c_{B}$, and the crash type $c_{T}$. To enable independent control over the influence of each signal during inference, we adopt a multi-condition extension of classifier-free guidance (CFG) from Equation 2, following formulations inspired by (liu2023compositionalvisualgenerationcomposable, ) and (pix2pix, ).
In standard CFG, both conditional and unconditional noise predictions are produced by a single model $\epsilon_{\theta}(\mathbf{x}_{t},c)$ and $\epsilon_{\theta}(\mathbf{x}_{t},\emptyset)$, respectively. However, as noted by recent work (uncond_priors, ), this setup can lead to poor unconditional priors and hinder conditional fidelity. To address this, we use two separate networks: the original pretrained base model $\epsilon_{\phi}$ to compute the unconditional predictions, and the fine-tuned Ctrl-Crash model $\epsilon_{\theta}$ to compute the conditional ones.
We define our multi-condition CFG formulation as follows:
$\hat{\epsilon}_{\theta,\phi}(\mathbf{x}_{t},c_{I},c_{B},c_{T})=\>$ $\epsilon_{\phi}(\mathbf{x}_{t},c_{I},\emptyset,\emptyset)$
@@ 3.5 Classifier-Free Guidance with Multi-Condition Factorization
$+\gamma_{T}\left[\epsilon_{\theta}(\mathbf{x}_{t},c_{I},c_{B},c_{T})-\epsilon_{\theta}(\mathbf{x}_{t},c_{I},c_{B},\emptyset)\right]$ (3)
Here, the first term $\epsilon_{\phi}(\mathbf{x}_{t},c_{I},\emptyset,\emptyset)$ represents the unconditional noise prediction, grounded on the initial image $c_{I}$ alone. We use the pretrained Stable Video Diffusion (SVD) model for this term, rather than computing a fully unconditional prediction (i.e., with $c_{I}$ removed), as we found the latter introduces substantial inference cost and minimal performance gain. The remaining terms use the fine-tuned model $\epsilon_{\theta}$ to compute conditional noise estimates. The scalar coefficients $\gamma_{B}$ and $\gamma_{T}$ control the relative guidance strength for the bounding box and crash type conditions, respectively. This factorized formulation allows us to modulate each conditioning channel independently at inference time, enabling fine-grained and interpretable control over the generated video’s spatial and semantic properties.
@@ 4.1 Quantitative Evaluation
We evaluate the generation quality of Ctrl-Crash across two main settings: general crash video generation in Table 1 and the impact of varying control signal strength in the prediction and reconstruction tasks in Table 2. All generated videos are 25 frames long at a resolution of $512\times 320$, and metrics are computed over 200 samples unless otherwise stated.
We evaluate generation quality using several video and frame-level metrics. Fréchet Video Distance (FVD) (unterthiner2019fvd, ) measures the similarity between the distributions of generated and real videos by analyzing video embeddings in I3D space under the assumption of Gaussian distributions; lower FVD values indicate a closer resemblance to real data. JEDi  luo2025jedi is a novel metric designed as an alternative to FVD by addressing its limitations through relaxing the assumption of Gaussian-distributed video feature embeddings, modifying the embedding space, and enabling convergence with a smaller sample size. In addition to these distributional-based video metrics, we report LPIPS (Learned Perceptual Image Patch Similarity) (zhang2018lpips, ), SSIM (Structural Similarity Index), and PSNR (Peak Signal-to-Noise Ratio) in Table 2, which evaluate frame-level fidelity and perceptual closeness relative to ground truth. These metrics provide complementary insights into how well the generated videos preserve appearance, structure, and temporal dynamics.
Table 1: Comparison of accident video generation quality across diffusion-based methods. We report FVD and JEDi scores ($\downarrow$ lower is better). For fair comparison, we compute the ground truth distribution from a set of 500 randomly sampled MM-AU (fang2024abductive, ) validation videos, to accommodate for methods without GT alignment (e.g., AVD2) (see Appendix F for more details). Scores marked with * are taken directly from the original papers and may not be strictly comparable due to differences in evaluation setup.
@@
OAVD img + bbox + text 5238* -
DrivingGen img + bbox + text 978.0* -
@@
Ctrl-V img + all bboxes 517.1 0.2910
Ctrl-Crash (Ours) img + all bboxes + GT crash type 449.5 0.1219
As shown in Table 1, Ctrl-Crash achieves the best results across both FVD and JEDi, indicating strong alignment with real crash dynamics and superior video quality. The SVD Base model, which serves as the foundation for Ctrl-Crash, performs poorly, as it was not trained for driving or crash-related content. Ctrl-V, while similar in architecture, lacks crash-specific training data and semantic control, leading to notable quality degradation near the crash event. AVD2 performs moderately well but exhibits inconsistent visual quality and weaker temporal coherence compared to Ctrl-Crash, as confirmed by FDV and JEDi. These results highlight the importance of both targeted training and structured control for crash simulation.
@@
Ours 0 (none) 422.1 0.3155 0.3856 0.5188 16.57
Ours 3 375.7 0.2949 0.3594 0.5434 17.27
Ours 9 353.3 0.2160 0.3392 0.5614 17.83
Ours 25 (all) 323.9 0.1219 0.3113 0.5836 18.33
Table 3: Effect of crash type conditioning on crash video generation quality. We evaluate Ctrl-Crash on the Crash Counterfactuals task by varying the crash type conditioning and nothing else. For each case, 200 videos were generated using the same initial image and three bounding box frames as initial context, but with different desired crash types. By comparing with the results from using the GT crash type (first line, taken from Table 2), we observe that the generated video quality is slightly worse or on par in almost all cases, suggesting the model can generate plausible alternatives while remaining visually close to the ground-truth video and while maintaining good video quality.
[…]
Ours 3 GT crash type 375.7 0.2949
Ours 3 0 - no crash 400.9 0.4514
Ours 3 1 - ego-only 379.5 0.3091
Ours 3 2 - ego/vehicle 372.9 0.3001
Ours 3 3 - vehicle-only 398.1 0.3856
Ours 3 4 - vehicle/vehicle 383.4 0.3416
Effect of Bounding Box Conditioning
We also study the impact of varying the number of bounding box frames used as conditioning for Ctrl-Crash, in Table 2. Specifically, we compare Crash Prediction (partial bbox inputs) and Crash Reconstruction (full bbox inputs) over 25-frame outputs. As shown in Table 2, generation quality improves consistently with the number of provided bounding box frames. This trend is visible across both distributional metrics (FVD, JEDi) and frame-level scores (LPIPS, SSIM, PSNR), and supports the hypothesis that denser spatial constraints lead to easier prediction tasks and more stable outputs. The results validate that Ctrl-Crash gracefully interpolates between unconditional prediction and fully supervised reconstruction.
@@ 4.2 Qualitative Evaluation
(a) Visual preference comparison in generated videos.
(b) Realism preference comparison in physical appearance.
Figure 3: User study results comparing generated videos from AVD2, DrivingGen and Ctrl-Crash with 40 participants across 5 crash types: Participants exhibit a strong preference for Ctrl-Crash generated videos, citing superior visual quality and physical realism. See Appendix I for more details.
We conducted a brief user study with $n=40$ participants who were asked to rank $k=3$ videos (from Ctrl-Crash, AVD2, and DrivingGen) across 5 different crash scenarios each. The participants consist of students from our lab who are not associated with the project in any way. No further demographic data was collected. The users were asked to rank the 3 videos in each of the 5 questions by (a) physical plausibility and (b) visual fidelity from best to worst. The users were required to choose a best/medium/worst video in each question and (visual/physical) category. We used the non-parametric Friedman test (friedman1937use, ) to determine with $p\leq 0.01$ that there is a method that is consistenly ranked higher than the others. We further used the Nemenyi Post-Hoc analysis (nemenyi1963distribution, ) to find that with $p\leq 0.01$, our method consistenly outperforms both AVD2 and DrivingGen in both physical realism and visual fidelity (see Figure 3).
Figure 4: Qualitative results comparing AVD2, DrivingGen, Ctrl-V, and Ctrl-Crash. The crash generated by AVD2 is visually shaky, with scenes that often lack consistency. Driving-Gen also produces low-quality and choppy videos. While Ctrl-V achieves good visual quality, it fails to generate realistic crash events. In contrast, Ctrl-Crash outperforms all baselines in both visual fidelity and scene consistency, while accurately modeling crash dynamics. Additional video demonstrations are available on the project page, and in Appendix A.1.
@@ 5 Conclusion
In this work, we introduced Ctrl-Crash, a controllable video diffusion framework that generates realistic car crash scenarios from a single frame, achieves state-of-the-art performance among diffusion-based methods, and enables counterfactual reasoning by varying spatial and semantic control inputs. To support training and evaluation, we also developed a processing pipeline for extracting bounding boxes from crash videos and released curated, annotated versions of MM-AU, RussiaCrash, and BDD100k to facilitate future research in crash simulation and generative modeling.
Despite its strong performance, our approach has several limitations, which motivates future work in this direction. Counterfactual outcomes can be hard to generate when initial scene conditions conflict with the desired crash type. The model also relies heavily on bounding boxes, making it sensitive to tracking errors—especially in fully conditioned reconstruction. With no bounding boxes conditioning, motion direction can be ambiguous, and 2D boxes struggle to capture rotation or orientation, limiting realism in behaviors like spinouts—future work may explore 3D bounding boxes or richer trajectory representations to overcome this. We envision Ctrl-Crash as a foundational tool for advancing the development of controllable generative models in safety-critical autonomous driving research.
Acknowledgments
We thank Samsung, the IVADO and the Canada First Research Excellence Fund (CFREF) / Apogée Funds, the Canada CIFAR AI Chairs Program, and the NSERC Discovery Grants program for financial support. We also thank Mila - the Quebec AI Institute for compute resources.
@@ Appendix A Additional Results
A.1 Additional Qualitative Results
Figure 5: Other diffusion-based video models—including state-of-the-art approaches—consistently fail to generate plausible crash scenarios. Top row: Nvidia Cosmos-Predict1-7B-Text2World with the prompt "On a highway two cars collide at very fast speeds head-on" produces a highly implausible scene, where the car on the left suddenly starts to levitate from its rear-end and a cloud of smoke resembling an explosion appears, followed by disjointed fragments of torn metal emerging from the ground that transforms in a dark vehicle rushing towards the camera. Middle row: OpenAI Sora with the prompt "At an intersection two cars collide with each other at full speed resulting in a crash" generates a car that spins erratically, changes shape and direction, and produces visible artifacts—culminating in a pile of twisted metal and glass. Bottom row: Ctrl-V, conditioned on a sequence of bounding boxes representing a head-on collision, renders two extremely blurry cars approaching each other. One car vanishes entirely at the moment of impact, while the other gradually fades away. Despite partially capturing the intent of the prompts, all models fail to produce a physically plausible and coherent crash, even when multiple samples are generated.
Despite the seemingly high quality video generation results produced by the latest wave of generative models, they systematically fail to generate physically plausible crashes and tend break down, yielding highly implausible imagery. For example in Figure 5 we show frames from Nvidia Cosmos [1], OpenAI Sora [3], and Ctrl-V [30]. In contrast, through our approach of sourcing crash data, preprocessing and annotating sequences with bounding boxes, crash type conditioning and our stochastic conditioning approach, our model can much more reliably generate physically plausible crash imagery and collision dynamics.
Other recent work [15][27] focus on car crash video generation, but lack the visual quality for convincing car crashes. As shown in the main paper, we conducted a user survey that shows that Ctrl-Crash significantly outperforms these existing methods for visual fidelity and physical plausibility of generated car crashes. We show additional qualitative comparisons in Figure 6 and Figure 7, along with many samples generated from Ctrl-Crash in Figure 8, Figure 9, and Figure 10.
Figure 6: Qualitative comparison of "rear-end crashes" between different methods. For each method, we show 5 frames from the video along with either a green check mark if there appears to have a crash in the video otherwise a red ’X’. From top to bottom: SVD (stable-video-diffusion-img2vid) [2] prompted with the initial frame from a rear-end crash video, we see some normal driving but very inconsistent lighting and color shades with visible distorsions. AVD2 [27], we see what appears to be a rear-end crash with a very distorted leading vehicle and background. DrivingGen [15], we see a rear-end crash with a leading vehicle that changes appearance every frame. Overall the video is very choppy with little temporal consistency. Cosmos (Cosmos-Predict1-7B-Video2World) [1], prompted with text suggesting a rear-end crash and 9 initial images where a car is rapidly approaching a truck, the predicted frames show the car unrealistically shrinking as it approaches the truck without any signs of a collision. Ctrl-V [30], prompted with a sequence of bounding-boxes suggesting a rear-end crash with a leading car, we see the leading car keep its distance and not crash occurs. Ctrl-Crash (ours): prompted with the same bounding box sequence as Ctrl-V and with the discrete crash type "ego/vehicle crash", we see a physically plausible rear-end collision with the ego vehicle visibly shaking from the impact. Visit our project page for animated video examples: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/
Figure 7: Qualitative comparison of "t-bone crashes" between different methods. For each method, we show 6 frames from the video along with either a green check mark if there appears to have a crash in the video otherwise a red ’X’. From top to bottom: SVD (stable-video-diffusion-img2vid) [2] prompted with the initial frame from a t-bone car crash video, we see blurry vehicle a distorted motion blur as it drives in front of the ego vehicle without any collision. AVD2 [27], we can make out what seems to be a t-bone crash with a heavily distorted black car. There are many artifacts and temporal inconsistencies which makes the sequence of events hard to follow. DrivingGen [15], a gray sedan drive in front of the ego vehicle progressively getting closer until it seems to collide with it. Motion is jerky and uneven between timesteps and the appearance of the gray car shapes almost every frame. Cosmos (Cosmos-Predict1-5B-Video2World) [1], prompted with creating a t-bone crash and 9 initial frames showing a car turn in front of the ego car, we see the leading car start to distort as the ego approaches it and then it shrivels and shrinks until it nearly disappears. Ctrl-V [30], prompted with a sequence of bounding-boxes suggesting a t-bone crash with a car incoming from the left, we see a car drive in from the left and then just passed the ego car without any collision. Ctrl-Crash (ours): prompted with the same bounding box sequence as Ctrl-V and with the discrete crash type "ego/vehicle crash", we see a physically plausible t-bone collision with the a black sedan incoming from the left. Visit our project page for animated video examples: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/
Figure 8: Ctrl-Crash qualitative results conditioned on an initial 9 bounding box frames (i.e., the first two frames of these sequences were conditioned on bounding box frames, but not the others). The animated videos for the examples presented above are provided in the supplementary material ZIP file and more video examples can be viewed at our project page: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/
Figure 9: Ctrl-Crash qualitative results conditioned on 25 (all) bounding box frames. The animated videos for the examples presented above are provided in the supplementary material ZIP file and more video examples can be viewed at our project page: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/
A.2 Additional Quantitative Results: Ablations

### v2

@@ 3.3 Conditioning Signals
Ctrl-Crash generates videos from three complementary conditioning signals that guide both the visual realism and semantic plausibility of the crash scenario (as illustrated in Figure 2):
Initial Frame (Scene Context): A single starting image grounds the generation process, providing the model with the appearance, layout, and environment of the driving scene before the crash occurs. It is encoded using a pretrained visual encoder (VAE) and fed into the main diffusion model as context.
Bounding-Box Trajectories (Spatial Control): The motion of road users is guided by sequences of 2D bounding boxes, which encode object positions, dimensions, unique track ID (via fill color), and classes (via border color) over time. These control frames are encoded using the same pretrained VAE as the initial image and processed by a ControlNet and injected into the denoising process to allow the model to follow specific trajectories and react to interactions in a physically plausible way. See E for a visualization.
Crash Type Label (Semantic Control): Crash types are represented by a discrete class label from five categories: (0) none, (1) ego-only, (2) ego/vehicle, (3) vehicle-only, and (4) vehicle/vehicle. These indicate which agents are involved in the crash, with, for example, “vehicle/vehicle” describing a crash between two non-ego agents (ego being the vehicle from which we observe the scene from). The crash type index is embedded and projected into the cross-attention layers of the ControlNet encoder, allowing the model to generate outcomes consistent with high-level semantic intent.
@@ 3.4 Training Pipeline
Our training strategy is divided into two phases. In the first stage, we fine-tune the SVD model on crash-heavy video data from our preprocessed version of the MM-AU dataset (see Section 3.2). The setup is an image-to-video task with an MSE loss in latent space (Equation 1). This improves the model’s ability to generate plausible driving videos, including crashes, though without explicit controllability. In the second stage, we freeze the fine-tuned base model and train a ControlNet adapter module with the same objective to introduce conditional controllability. This stage allows us to direct the generation using spatial and semantic control signals—namely, bounding box frames and crash type identifiers. Additional training and implementation details are provided in G for reproducibility.
Conditioning Masking. To improve robustness and support flexible guidance at inference, we apply randomized masking of control signals during training. For bounding boxes, we use a temporal dropout strategy: a random timestep $k$ is sampled, and all subsequent frames are replaced with a learnable null embedding to avoid misinterpreting missing input as the absence of agents. A curriculum schedule gradually increases the masking ratio to encourage generalization to partial control. For semantic inputs (initial image and crash type), we apply joint masking with a fixed probability schedule: with 10% chance only the image is masked, 10% only the crash type, and 10% both—encouraging balanced reliance across modalities and enabling effective classifier-free guidance.
3.5 Multi-Condition Classifier-Free Guidance
Ctrl-Crash supports three conditioning modalities: initial image $c_{I}$, bounding box frames $c_{B}$, and crash type $c_{T}$. To enable independent control over each during inference, we extend classifier-free guidance (CFG) following [29, 17], and denote the final noise estimate under multi-condition CFG as $\hat{\epsilon}_{\theta,\phi}$:
$\hat{\epsilon}_{\theta,\phi}(\mathbf{x}_{t},c_{I},c_{B},c_{T})=$ $\epsilon_{\phi}(\mathbf{x}_{t},c_{I},\emptyset,\emptyset)$ (3)
@@ 3.5 Multi-Condition Classifier-Free Guidance
$\!+\gamma_{T}\left[\epsilon_{\theta}(\mathbf{x}_{t},c_{I},c_{B},c_{T})-\epsilon_{\theta}(\mathbf{x}_{t},c_{I},c_{B},\emptyset)\right]$
Unlike standard CFG (Equation 2), which uses the same model for both conditional and unconditional predictions, we follow [30] and decouple these roles: $\epsilon_{\phi}$ (the off-the-shelf SVD model) predicts the unconditional term, while $\epsilon_{\theta}$ (the fine-tuned Ctrl-Crash model) handles conditional terms. This avoids degraded unconditional priors and improves conditional fidelity. The coefficients $\gamma_{B}$ and $\gamma_{T}$ modulate the strength of bounding box and crash type conditionings, respectively, enabling flexible and interpretable control at inference.
@@ 4.1 Quantitative Evaluation
Table 1: Comparison of accident video generation quality across diffusion-based methods. We report FVD and JEDi scores ($\downarrow$ lower is better). Scores marked with * are taken directly from the original papers and may not be strictly comparable due to differences in evaluation setup. The ”Conditions” column describes inputs used as control signals. ”img”: initial image frame, ”BB”: bounding box frames, ”text”: textual description, ”action”: discrete crash category value. See Figure 4 for visuals.
@@
OAVD img + BB + text 5238* -
DrivingGen img + BB + text 978.0* -
@@
Ctrl-V img + BB 517.1 0.2910
Ctrl-Crash img + BB + action 449.5 0.1219
We evaluate the generation quality of Ctrl-Crash across two core settings: (1) general crash video generation in Table 1; and (2) varying the number of bounding box guidance frames provided in the prediction task to assess the impact of partial trajectory information in Table 2. All generated videos are 25 frames long at a resolution of $512\times 320$, and metrics are computed over 200 samples unless otherwise specified.
Figure 3: Counterfactual Crash Generation: this diagram demonstrates the ability of our model to generate counterfactual crashes (Middle: no crash, Bottom: ego/car crash) while beginning from the initial frame and 3 bounding-boxes frames of the real video (Top: the real car crash).
We evaluate generation quality using several video and frame-level metrics. Fréchet Video Distance (FVD) [31] measures the similarity between the distributions of generated and real videos by analyzing video embeddings in I3D space under the assumption of Gaussian distributions; lower FVD values indicate a closer resemblance to real data. JEDi  [32] is a novel metric designed as an alternative to FVD by addressing its limitations through relaxing the assumption of Gaussian-distributed video feature embeddings, modifying the embedding space, and enabling convergence with a smaller sample size. In addition to these distributional-based video metrics, we report LPIPS (Learned Perceptual Image Patch Similarity) [33], SSIM (Structural Similarity Index), and PSNR (Peak Signal-to-Noise Ratio) in Table 2, which evaluate frame-level fidelity and perceptual closeness relative to ground truth, averaged over the video frames. These metrics provide complementary insights into how well the generated videos preserve appearance, structure, and temporal dynamics.
As shown in Table 1, Ctrl-Crash achieves the best results across both FVD and JEDi among compared methods, indicating strong alignment with real crash dynamics and superior video quality. For fair comparison, we compute the ground truth (GT) distribution for FVD and JEDi from a set of 500 randomly sampled MM-AU [11] validation videos, to accommodate for methods without GT alignment (e.g., AVD2) (see Appendix F in supplementary material for more details). The SVD Base model, which serves as the foundation for Ctrl-Crash, performs poorly, as it was not trained for driving or crash-related content. Ctrl-V, while similar in architecture, lacks crash-specific training data and semantic control, leading to notable quality degradation near the crash event. AVD2 performs moderately well but exhibits inconsistent visual quality and weaker temporal coherence compared to Ctrl-Crash, as confirmed by FVD and JEDi. These results highlight the importance of both targeted training and structured control for crash simulation.
@@
0 422.1 0.3155 0.3856 0.5188 16.57
3 375.7 0.2949 0.3594 0.5434 17.27
9 353.3 0.2160 0.3392 0.5614 17.83
25 323.9 0.1219 0.3113 0.5836 18.33
We also study the impact of varying the number of bounding box frames used as conditioning for Ctrl-Crash, in Table 2. Specifically, we compare using partial bounding-box frame conditioning from zero frames to a few bounding box frames (Crash Prediction task) all the way to a fully defined motion sequence by providing all 25 frames (Crash Reconstruction task). As shown in Table 2, generation quality improves consistently with the number of provided bounding box frames. This trend is visible across both distributional metrics (FVD and JEDi) and frame-level scores (LPIPS, SSIM, PSNR), and supports the hypothesis that denser spatial constraints lead to easier prediction tasks and more stable outputs. The results validate that Ctrl-Crash gracefully interpolates between unconditional prediction and fully supervised reconstruction.
@@ 4.2 Qualitative Evaluation
Figure 4: Qualitative comparison of AVD2, DrivingGen, Ctrl-V, and Ctrl-Crash. Ctrl-Crash achieves superior visual fidelity and realistic crash dynamics, outperforming baselines that suffer from low quality, inconsistency, or lack of plausible crash events.
We qualitatively tested leading video diffusion models—OpenAI’s Sora [34], Nvidia’s Cosmos [19], and DeepMind’s Veo3 [35]—for their ability to generate car crash scenarios. While these models produce high-resolution, cinematic visuals, they consistently struggle to depict physically plausible accidents. In our experiments, Cosmos and Sora failed to generate convincing crashes across several prompt variations, while Veo3, which we tested with a single sample, showed limited realism despite momentarily resembling a crash (see A.2 for exact prompts). These examples, shown in Figure 5, suggest that current general-purpose models lack the exposure, the fine-grained control, and physical grounding required for realistic accident simulation.
Figure 5: State-of-the-art text-to-video diffusion models struggle to generate a realistic car accident. Top: Nvidia Cosmos (7B Text2World), Middle: OpenAI Sora, Bottom: DeepMind Veo3.
Additionally, we conducted a brief user study to evaluate visual quality and physical realism of the generated videos based on human preference. Results show clear and significant preference for Ctrl-Crash among similar methods. Results and details are presented in I.
@@ 5 Conclusion
In this work, we introduced Ctrl-Crash, a controllable video diffusion framework that generates realistic car crash scenarios from a single frame, which achieves state-of-the-art performance among diffusion-based methods designed for crashes, and enables counterfactual reasoning by varying spatial and semantic control inputs. To support training and evaluation, we also developed a processing pipeline for extracting bounding boxes from crash videos and released curated, annotated versions of MM-AU, RussiaCrash, and BDD100k to facilitate future research in crash simulation and generative modeling.
Despite its strong performance, our approach has several limitations, motivating promising directions for future work. Counterfactual outcomes can be difficult to generate when the initial scene strongly suggests a different crash outcome than the one being conditioned on. The model also relies heavily on 2D bounding boxes, which makes it sensitive to tracking errors—particularly in fully conditioned reconstruction. Moreover, 2D boxes often lack information about agent orientation and rotation, limiting realism for behaviors such as spinouts or rollovers. Future extensions may explore 3D bounding boxes or richer trajectory representations to overcome this. Another limitation is the use of a fixed set of five discrete crash type labels to define semantic intent. While simple and effective, this formulation constrains expressiveness. Incorporating natural language as a conditioning signal could unlock more nuanced control and allow the model to interpret detailed scene descriptions or intentions that go beyond predefined categories and bounding-box information. We envision Ctrl-Crash as a foundation for building more general, interpretable, and controllable generative models for safety-critical autonomous driving research.
Acknowledgments: We thank Samsung, the IVADO and the Canada First Research Excellence Fund (CFREF) / Apogée Funds, the Canada CIFAR AI Chairs Program, and the NSERC Discovery Grants program for financial support. We also thank Mila - the Quebec AI Institute for compute resources.
@@ Appendix A Additional Results
A.1 Additional Quantitative Results: Ablations